# Pangu DeepDiver: Adaptive Search Intensity Scaling via Open-Web Reinforcement Learning

### v1

@@ Abstract
Information seeking demands iterative evidence gathering and reflective reasoning, yet large language models (LLMs) still struggle with it in open-web question answering. Existing methods rely on static prompting rules or training with Wikipedia-based corpora and retrieval environments, limiting adaptability to the real-world web environment where ambiguity, conflicting evidence, and noise are prevalent. These constrained training settings hinder LLMs from learning to dynamically decide when and where to search, and how to adjust search depth and frequency based on informational demands. We define this missing capacity as Search Intensity Scaling (SIS)–the emergent skill to intensify search efforts under ambiguous or conflicting conditions, rather than settling on overconfident, under-verification answers.
To study SIS, we introduce WebPuzzle, the first dataset designed to foster information-seeking behavior in open-world internet environments. WebPuzzle consists of 24K training instances and 275 test questions spanning both wiki-based and open-web queries. Building on this dataset, we propose DeepDiver, a Reinforcement Learning (RL) framework that promotes SIS by encouraging adaptive search policies through exploration under a real-world open-web environment. Experimental results show that Pangu-7B-Reasoner empowered by DeepDiver achieve performance on real-web tasks comparable to the 671B-parameter DeepSeek-R1. We detail DeepDiver’s training curriculum from cold-start supervised fine-tuning to a carefully designed RL phase, and present that its capability of SIS generalizes from closed-form QA to open-ended tasks such as long-form writing. Our contributions advance adaptive information seeking in LLMs and provide a valuable benchmark and dataset for future research.
@@ 1 Introduction
Information seeking (Wilson, 1999) is a fundamental cognitive skill that involves evidence gathering, reflective reasoning, and the resolution of conflicting information. Despite significant advancements in artificial intelligence, LLMs continue to struggle with replicating such information-seeking behaviors. Knowledge-intensive question answering, a central challenge for LLMs, requires a robust capability for information seeking. Current models often fail to determine when and what information to seek, verify the relevance of evidence, and reason effectively over noisy or conflicting contexts.
@@
Iterative Retrieval-Augmented Generation (RAG) (Lewis et al., 2021) frameworks have been proposed to address these challenges by alternating between retrieval and reasoning. Existing approaches generally fall into two categories: prompting-based and task-specific supervised fine-tuning (SFT). Prompting-based methods leverage predefined rules or in-context learning (ICL) (Brown et al., 2020), forcing the LLM to follow a specific pipeline that iteratively searches and reasons to complete complex tasks (Jiang et al., 2023; Shao et al., 2023; Trivedi et al., 2023; Yue et al., 2025; Li et al., 2025). However, the fixed reasoning and searching strategies provided in the prompts limit their adaptability to complex, dynamic problems. In contrast, supervised fine-tuning methods train models to improve retrieval and reasoning capabilities (Asai et al., 2023; Yu et al., 2024), generally yielding better performance. However, these methods often internalize inference patterns tied to the training corpus, restricting generalization to more dynamic or unseen situations.
@@
Recently, reinforcement learning (RL) (Sutton and Barto, 2018; Kaelbling et al., 1996) has been applied to enhance inference-time reasoning in LLMs, enabling iterative refinement and exploration of reasoning (DeepSeek-AI et al., 2025; Team, 2025; OpenAI, 2024a). Several studies have integrated RL into iterative RAG frameworks, encouraging models to explore diverse reasoning paths and rewarding accurate outcomes (Jin et al., 2025; Song et al., 2025; Chen et al., 2025a; Zheng et al., 2025). However, these works predominantly train and evaluate their methods on well-structured datasets such as HotpotQA (Yang et al., 2018), which are based on corpora like Wikipedia. In such settings, many tasks can be effectively solved using solely the LLMs’ internal knowledge, and the introduced search environments are “clean,” containing minimal noise or conflicting information. In contrast, real-world search environments are inherently more complex, characterized by noisy, inconsistent, and unreliable sources. This discrepancy limits the generalizability of the reported “incentivized search capabilities” to more realistic, open-ended information-seeking scenarios.
@@
Consequently, due to flawed training data and simplistic search environments, previous works fail to discover a critical capability in LLMs that we define as Search Intensity Scaling (SIS): the ability to dynamically scale search frequency and depth as information demands increase, rather than settling for overconfident, under-verified answers. This ability enables LLMs to tackle more complex, information-intensive problems under open-web environment, where ambiguity, conflicting evidence, and noise are prevalent. To foster this emergent capability, we introduce WebPuzzle, a dataset designed to evaluate information-seeking capabilities in real-world search environments. WebPuzzle contains 24k training samples and 275 human-annotated test examples, covering tasks solvable with Wikipedia content as well as broader open-domain queries extracted from open-web environment. Even Wikipedia subsets are rigorously validated to require external retrieval, ensuring a realistic assessment of LLMs’ information seeking behaviors. Along with WebPuzzle, we introduce DeepDiver, an RL-driven search and reasoning framework trained on this dataset. DeepDiver follows a conventional cold-start SFT then RL pipeline, enhanced with carefully designed tool-calling reward assignments and scheduling mechanisms. Critically, DeepDiver interacts with real-world search engines throughout the training phase. This interaction enables it to continuously refine and denoise retrieved documents, forcing the LLM locate scattered yet salient information within noisy open-web environments—ultimately leading to more accurate answers.
@@
Through systematic empirical analysis, we identify critical factors that influence model behavior, including the search intensity, the training environment, and the generalization capabilities. Our analysis reveals several key insights: (1) DeepDiver exhibits exceptional information-seeking ability via SIS, where the depth and frequency of searching are proportional to both problem difficulty and the model’s performance. (2) Compared to the “clean” Wiki-based environment, WebPuzzle and real-world open-web search environments better support complex reasoning beahviours, guiding LLMs to actively supplement evidence, resolve conflicts, verify content, and reflect for self-correction. (3) RL training significantly enhances the generalization capability of LLMs, enabling the transition from closed-ended to open-ended problems. In conclusion, this report underscores the potential of our proposed WebPuzzle and DeepDiver to foster emergent adaptive search behaviors— search intensity scaling—in LLMs. This significantly enhances LLM’s ability to perform adaptive, verifiable, and scalable information seeking, providing a promising direction for future advancements in knowledge-intensive problem solving.
@@ 2.1 Iterative RAG
At each iteration $t\in\{1,2,\dots,T\}$, the model maintains a reasoning history:
$\mathcal{H}_{t-1}=\{q,(r_{1},s_{1},d_{1}),\dots,(r_{t-1},s_{t-1},d_{t-1})\}$ (1)
where $r_{i}$ represents the intermediate reasoning CoT generated at round $i$, $s_{i}$ denotes search queries, and $d_{i}$ denotes retrieved documents from web search.
At each round $t$, conditioned on history $\mathcal{H}_{t-1}$, the model first generates intermediate reasoning $r_{t}$ to analyze the current status:
$r_{t}\sim p(r_{t}\mid\mathcal{H}_{t-1})$ (2)
Then, based on the reasoning $r_{t}$, the model selects one of two actions:
  • Search: Generate additional queries $s_{t}$ and retrieve supporting documents $d_{t}$:
    $s_{t}\sim p(s_{t}\mid\mathcal{H}_{t-1},r_{t}),\quad d_{t}=\text{Retrieval}(s_{t})$ (3)
  • Answer: Finalize the answer $a$ to question $q$:
    $a\sim p(a\mid\mathcal{H}_{t-1},r_{t})$ (4)
This iterative reasoning and retrieval process continues until the model chooses the answer action, resulting in a final answer that is well-supported by retrieved evidence and explicit reasoning steps.
@@ 2.2 Information Seeking Behaviour
Figure 1: Illustration of four key information-seeking behaviors: (a) Evidence Gathering & Supplements, where the model formulates queries to fill knowledge gaps by retrieving supporting documents; (b) Conflict Resolution, where the model identifies and resolves contradictions in retrieved information; (c) Verification & Denoising, which involves cross-checking facts and filtering out noisy or irrelevant data; and (d) Reflection & Correction, where the model revisits earlier reasoning steps to correct mistakes and refine its understanding.
@@
We define information seeking behaviour within iterative RAG frameworks as a structured decision-making process, where at each iteration the model adopts specific strategies to resolve uncertainties, improve evidence quality, and enhance the overall reliability of answers. Formally, at iteration $t$, conditioned on the reasoning history $\mathcal{H}_{t-1}$ and current intermediate reasoning $r_{t}$, the model exhibits several strategies to guide its search and reasoning processes.
Inspired by the findings of  Gandhi et al. (2025), we categorize these strategies into four types of information seeking behaviours: Evidence Gathering & Supplements, Conflict Resolution, Verification & Denoising, and Reflection & Correction. Each behaviour represents a specific mode of interaction with the retrieval environment, addressing real-world, open-domain queries:
Evidence Gathering & Supplements.
The model actively seeks to fill identified knowledge gaps by formulating targeted queries $s_{t}$ and retrieving supporting documents $d_{t}$. Formally:
$(s_{t},d_{t})\sim p(s_{t},d_{t}\mid\mathcal{H}_{t-1},r_{t}),\quad\text{where }d_{t}=\text{Retrieval}(s_{t}).$ (5)
This strategy is exemplified by traditional question-answering datasets such as 2Wiki (Ho et al., 2020), HotpotQA (Yang et al., 2018), and FRAMES (Krishna et al., 2024), primarily relying on structured knowledge sources (e.g., Wikipedia).
Conflict Resolution.
In scenarios where retrieved information contains contradictions or discrepancies, the model explicitly reasons about inconsistencies and evaluates competing claims to resolve conflicts.
Verification & Denoising.
Given noisy or irrelevant retrieved information, the model cross-checks facts and isolates trustworthy information, thereby performing verification and denoising.
Reflection & Correction.
The model periodically reassesses its reasoning trajectory, revisits earlier assumptions, and explicitly corrects or refines previous reasoning steps, enabling iterative refinement and adaptive reasoning.
Formally, these three behaviours (Conflict Resolution, Verification & Denoising, and Reflection & Correction) can be represented generally as generating reasoning steps $r_{t}$ conditioned on the accumulated reasoning history and retrieved documents:
$r_{t}\sim p(r_{t}\mid\mathcal{H}_{t-1},d_{t}),$ (6)
where specific conditions (e.g., presence of contradiction, noise, or previous mistakes) differ according to each behaviour.
@@
Existing works adopting the wiki-based datasets, limiting their scope to structured and well-organized knowledge bases, and thus predominantly emphasize “Evidence Gathering & Supplements”. To prove this observation, we show a detailed analysis in Section 5.6. In contrast, our proposed WebPuzzle and the real-world searching environment explicitly necessitate employing all four behaviours, thereby reflecting a more comprehensive and realistic scenario for real-world problem-solving with web-searching. More details about WebPuzzle will be included in Section 3.1
2.3 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is a reinforcement learning algorithm designed to enhance the efficiency of Proximal Policy Optimization (PPO) (Schulman et al., 2017) by eliminating the need for a critic network. Specifically, GRPO samples multiple outputs from a previous policy for a given prompt and computes their average reward to serve as a dynamic baseline. The advantage of each output is defined relative to this baseline, resulting in positive advantages for outputs exceeding the baseline and negative advantages otherwise. Formally, the GRPO objective is defined as follows:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}\left[\sum_{i=1}^{G}\min\left(\frac{\pi_{\theta}(o_{i})}{\pi_{\theta_{\text{old}}}(o_{i})}A_{i},\;\text{clip}\left(\frac{\pi_{\theta}(o_{i})}{\pi_{\theta_{\text{old}}}(o_{i})},1-\epsilon,1+\epsilon\right)A_{i}\right)-\beta\,\mathbb{D}_{\text{KL}}\left(\pi_{\theta}\,\|\,\pi_{\text{ref}}\right)\right],$ (7)
where the relative advantage $A_{i}$ is computed as:
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\dots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\dots,r_{G}\})}.$ (8)
GRPO retains PPO’s clipping strategy and incorporates a KL-divergence term for regularization to ensure the stable training of RL. By utilizing this relative scoring scheme, GRPO obviates the requirement of a separate value estimation network, thereby simplifying the training procedure and reducing computational overhead.
@@ 3 Method
In this section, we discuss the details of our approach. We begin by introducing WebPuzzle, a dataset designed to address real-world reasoning and search challenges. Next, we describe DeepDiver, a RL-based training framework aimed at enhancing LLMs with robust capabilities introduced in Section  2.2.
@@ 3.1 WebPuzzle
Unlike existing open-domain QA datasets based on Wikipedia (Yang et al., 2018; Press et al., 2023; Krishna et al., 2024) where LLMs often perform well using only internal knowledge, we introduce WebPuzzle, a dataset designed to evaluate LLMs’ ability to locate and reason over noisy, scattered information on the open web. Figure LABEL:webpuzzle_pipeline illustrates our data synthesis and curation processes.
Candidate Data Generation We collect candidate data from Wiki-corpus and real-user queries with retrieved webpages from our deployed smart assistant service. Our generation involves two approaches: (1) Cross-page question generation, where an LLM extracts facts from web pages to generate “inverted” questions, answers and checklists (Tang and Yang, 2024; Krishna et al., 2024)—applied only to open web pages as Wiki-corpus tends to produce overly-simple questions; and (2) Riddle creation, where the LLM selects distinctive entity attributes and applies obfuscation or generalization to create challenging problems, with original entities as labels. Two examples are shown below.
{CJK*}
UTF8bsmi
Example of Cross-page Question
2024年10月31日,一名身高125cm的普通成人游客购买深圳欢乐谷万圣节夜场票,需支付多少钱?
Translation: On October 31st, 2024, how much does an ordinary adult tourist with a height of 125 cm need to pay for the night ticket of the Halloween event at Happy Valley Shenzhen?
Solution: 149 RMB
{CJK*}
UTF8gbsn
Example of Riddle
{CJK*} UTF8gbsn
某中国社交媒体应用因美国政府的禁令而意外获得大量关注,用户自称为“难民”并迁移到该平台。该应用在苹果应用商店一度下载量排名第一,超过ChatGPT。其母公司总部位于上海。这是哪个应用?
Translation: A Chinese social media app unexpectedly gained a great deal of attention due to a U.S. government ban. Users refer to themselves as “refugees” and have migrated to this platform. At one point, the app ranked first in downloads on the Apple App Store, surpassing ChatGPT. Its parent company is headquartered in Shanghai. Which app is this?
Solution: 小红书 / Rednote / Xiaohongshu
footnotetext: The WebPuzzle dataset will be released after privacy and safety review.
Difficulty Assessment To ensure stable RL training with consistent reward signals, we tag each problem’s difficulty level, enabling a data mixture strategy that prevents all-zero rewards, which could lead to training collapse. For each problem, we test DeepSeek-R1 four times, using the number of correct answers to determine difficulty. Formally, let $N_{\text{correct}}$ denote the number of correct answers out of the 4 tests. The difficulty level $D$ of a problem can be determined as:
$D=\begin{cases}\text{easy}&\text{if }N_{\text{correct}}=4\\ \text{medium}&\text{if }N_{\text{correct}}=2\text{ or }3\\ \text{hard}&\text{if }N_{\text{correct}}=1\\ \text{outlier}&\text{if }N_{\text{correct}}=0\end{cases}$ (9)
The statistics of the dataset are presented in Table 1, we detail the tagging workflow in Appendix C.3.
Data Category Training Data Num Evaluation Data Num
Easy Medium Hard Outliers ALL Easy Medium Hard Outliers ALL
Cross-Page QA 2553 2451 1404 3970 4500 5 54 27 44 130
Open&Wiki Riddle 5566 2956 1409 3375 2500 0 37 33 75 145
Total Set 8119 5407 2813 7345 23684 5 91 60 119 275
Test Set Annotation Unlike the training set which used LLM labeling, our test set was manually annotated by five human experts using an open-web search engine. From 500 seed samples, experts followed the principles in Appendix B.1 to ensure meaningful evaluation of LLMs’ information-seeking behaviors. Through iterative annotation, we finalized 275 samples for testing.

### v2

# DeepDiver: Adaptive Web-Search Intensity Scaling via Reinforcement Learning

@@ Abstract
Information seeking demands iterative evidence gathering and reflective reasoning, yet large language models (LLMs) still struggle with it in open-web question answering. Existing prompting and supervised fine-tuning (SFT) methods remain fixed by prompt rules or training corpora, and are usually benchmarked only on well-structured wiki sources, limiting real-world adaptability. We introduce WebPuzzle, a 24k-sample training and 275-sample test benchmark that evaluates information seeking on the live internet, across both wiki and open-domain queries. Leveraging 7k WebPuzzle instances, we develop DeepDiver, a reinforcement-learning (RL) framework that cultivates Search Intensity Scaling (SIS)—an emergent ability to escalate search frequency and depth instead of settling on overconfident, under-evidenced answers. With SIS, Qwen2.5-7B-Instruct and Pangu-7B-Reasoner attain performance on real-web tasks comparable to the 671B-parameter DeepSeek-R1. We detail DeepDiver’s curriculum from cold-start SFT to a well designed RL procedure, and show that its seeking policy generalized from closed-ended queries to open-ended generation such as long-form writing. Our results advance adaptive information seeking in LLMs and provide a rigorous benchmark for future work.
@@ 1 Introduction
Information seeking [wilson1999models] is a fundamental cognitive skill that involves iterative evidence gathering, reflective reasoning, and the resolution of conflicting information. Despite significant advancements in artificial intelligence, LLMs continue to struggle with replicating such information-seeking behaviors. Knowledge-intensive question answering, a central challenge for LLMs, requires a robust capability for information seeking. Current models often fail to determine when and what information to seek, verify the relevance of evidence, and reason effectively over noisy or conflicting contexts.
@@
Iterative Retrieval-Augmented Generation (RAG) [lewis2021retrievalaugmentedgenerationknowledgeintensivenlp] frameworks have been proposed to address these challenges by alternating between retrieval and reasoning. Existing approaches generally fall into two categories: prompting-based and task-specific supervised fine-tuning (SFT). Prompting-based methods leverage predefined rules or in-context learning (ICL) [brown2020languagemodelsfewshotlearners], forcing the LLM to follow a specific pipeline to complete complex tasks [jiang2023activeretrievalaugmentedgeneration, shao2023enhancingretrievalaugmentedlargelanguage, trivedi2023interleavingretrievalchainofthoughtreasoning, yue2025inferencescalinglongcontextretrieval, li2025searcho1agenticsearchenhancedlarge]. However, the fixed reasoning flow provided in the prompts limits their adaptability to complex, dynamic problems. In contrast, supervised fine-tuning methods train models to improve retrieval and reasoning capabilities [asai2023selfraglearningretrievegenerate, yu2024autoragautonomousretrievalaugmentedgeneration], generally yielding better performance. However, these methods often internalize inference patterns tied to the training corpus, restricting generalization to more dynamic or unseen situations.
@@
Recently, reinforcement learning (RL) [Sutton1998, kaelbling1996reinforcementlearningsurvey] has been applied to enhance inference-time reasoning in LLMs, enabling iterative refinement and exploration of reasoning [deepseekai2025deepseekr1incentivizingreasoningcapability, qwq32b, openai2024learning]. Several studies have integrated RL into iterative RAG frameworks, encouraging models to explore diverse reasoning paths and rewarding accurate outcomes [jin2025search, song2025r1searcherincentivizingsearchcapability, chen2025research, zheng2025deepresearcher]. However, these works predominantly train and evaluate their methods on well-structured datasets such as HotpotQA [yang2018hotpotqadatasetdiverseexplainable], which are based on corpora like Wikipedia. In such settings, many tasks can be effectively solved using the LLMs’ internal knowledge, and the introduced search environments are “clean,” containing minimal noise or conflicting information. In contrast, real-world search environments are inherently more complex—characterized by noisy, inconsistent, and unreliable sources. This discrepancy limits the generalizability of the reported “incentivized search capabilities” to more realistic, open-ended information-seeking scenarios.
@@
To investigate RL-guided LLM behaviors in more realistic, open-domain scenarios, we introduce WebPuzzle, a dataset designed to evaluate information-seeking capabilities in real-world search environments. WebPuzzle contains 24k training samples and 275 human-annotated test examples, covering tasks solvable with Wikipedia content as well as broader open-domain queries extracted from open-web environment. Even Wikipedia subset are rigorously validated to require external retrieval, ensuring a realistic assessment of LLMs’ search behaviors. Along with WebPuzzle, we introduce DeepDiver, an RL-driven search and reasoning framework trained on this dataset. DeepDiver interacts with real-world search engines, continuously refining and denoising retrieved documents to provide accurate answers. A key innovation of DeepDiver is the emergent capability of search intensity scaling (SIS), which dynamically scales up the search frequency and depth as information demands increase. This enables LLMs to tackle more complex, information-intensive problems under open-web environment. Together, WebPuzzle and DeepDiver provide a comprehensive framework for developing and examining information seeking ability of LLMs, offering a promising approach for knowledge-intensive problem solving.
@@
Through systematic empirical analysis, we identify critical factors that influence model behavior, including the search intensity, the training environment, and the generalization capabilities. Our analysis reveals several key insights: (1) DeepDiver exhibits exceptional information-seeking ability via adaptive SIS, where the depth and frequency of searching proportional to both problem difficulty and the model’s performance. (2) Compared to the “clean” Wiki-based environment, WebPuzzle and real-world search settings better support complex reasoning beahviours, guiding LLMs to actively supplement evidence, resolve conflicts, verify content, and reflect for self-correction. (3) RL training significantly enhances the generalization capability of LLMs, enabling the transition from closed-ended to open-ended problems. In conclusion, our method underscores the potential of reinforcement learning to foster emergent adaptive search behaviors—specifically, search intensity scaling—in LLMs. This significantly enhances their ability to perform adaptive, verifiable, and scalable information seeking, providing a promising direction for future advancements in knowledge-intensive problem solving.
@@ 2.1 Iterative RAG
At each iteration $t\in\{1,2,\dots,T\}$, the model maintains a reasoning history $\mathcal{H}_{t-1}=\{q,\,(r_{1},s_{1},d_{1}),\dots,(r_{t-1},s_{t-1},d_{t-1})\}$, where $r_{i}$ represents the intermediate CoT generated at round $i$, $s_{i}$ denotes search queries, and $d_{i}$ denotes retrieved documents from web search.
At round $t$, conditioned on history $\mathcal{H}_{t-1}$, the model first generates intermediate reasoning $r_{t}\sim p\bigl(r_{t}\mid\mathcal{H}_{t-1}\bigr)$ to analyze the current status. Then, based on the reasoning $r_{t}$, the model selects one of two actions: (1) Search: generate additional queries $s_{t}\sim p\bigl(s_{t}\mid\mathcal{H}_{t-1},r_{t}\bigr)$ and retrieve supporting documents $d_{t}=\text{Retrieval}(s_{t})$; (2) Answer: finalize the answer $a\sim p\bigl(a\mid\mathcal{H}_{t-1},r_{t}\bigr)$ to question $q$. This iterative reasoning-and-retrieval process continues until the model chooses the answer action, resulting in a final answer that is well supported by retrieved evidence and explicit reasoning steps.
Figure 1: Illustration of four key information-seeking behaviors: (a) Evidence Gathering & Supplements (b) Conflict Resolution (c) Verification & Denoising and (d) Reflection & Correction.
@@ 2.2 Information Seeking Behaviour
We define information seeking behaviour within iterative RAG frameworks as a structured decision-making process: at each iteration the model adopts specific strategies to resolve uncertainties, improve evidence quality, and enhance the overall reliability of answers. Formally, at iteration $t$, conditioned on the reasoning history $\mathcal{H}_{t-1}$ and current intermediate reasoning $r_{t}$, the model exhibits several strategies to guide its search and reasoning processes.
Inspired by the findings of gandhi2025cognitivebehaviorsenableselfimproving, we categorize these strategies into four types of information seeking behaviours: (1) Evidence Gathering & Supplements, where the model actively seeks to fill identified knowledge gaps by formulating targeted queries $s_{t}$ and retrieving supporting documents $d_{t}$, formally represented as $(s_{t},d_{t})\sim p\bigl(s_{t},d_{t}\mid\mathcal{H}_{t-1},r_{t}\bigr),\;\text{where }d_{t}=\text{Retrieval}(s_{t}).$ This strategy is exemplified by traditional question-answering datasets such as 2Wiki [ho2020constructingmultihopqadataset], HotpotQA [yang2018hotpotqadatasetdiverseexplainable], and FRAMES [krishna2024factfetchreasonunified]; (2) Conflict Resolution, where the model reasons about inconsistencies and evaluates competing claims when retrieved information contains contradictions; (3) Verification & Denoising, where the model cross-checks facts and isolates trustworthy information from noisy or irrelevant retrieved content; and (4) Reflection & Correction, where the model periodically re-assesses its reasoning trajectory, revisits earlier assumptions, and explicitly corrects previous reasoning steps for iterative refinement. The latter three behaviours can be represented generally as generating reasoning steps $r_{t}\sim p\bigl(r_{t}\mid\mathcal{H}_{t-1},d_{t}\bigr),$ where the specific conditions (e.g. presence of contradiction, noise, or previous mistakes) differ according to each behaviour.
@@
Existing works adopting the wiki-based datasets, limiting their scope to structured and well-organized knowledge bases, and thus predominantly emphasize “Evidence Gathering & Supplements”. To prove this observation, we show an detailed analysis in Appendix A.2. In contrast, our proposed WebPuzzle and the real-world searching environment explicitly necessitates employing all four behaviours, thereby reflecting a more comprehensive and realistic scenario for real-world problem-solving with web-searching. More details about WebPuzzle will be included in section 3.1.
@@ 3 Method
In this section, we discuss the details of our approach. We begin by introducing WebPuzzle, a dataset designed to address real-world reasoning and search challenges. Next, we describe DeepDiver, a reinforcement learning-based training framework aimed at enhancing LLMs with robust capabilities introduced in section  2.2.
Figure 2: WebPuzzle pipeline. Above: Candidate Generation: Wiki and open-web pages yield QA pairs via (i) Cross-Page QA and (ii) Riddle pipelines, grouped as Cross-Page QA, Open Riddle, and Wiki Riddle. Below: Difficulty Tagging: Each sample is tagged (easy/medium/hard) for adaptive mixing in RL; DeepDiver is trained on a curated 7k-sample mix.
@@ 3.1 WebPuzzle
Unlike existing open-domain QA datasets based on Wikipedia [yang2018hotpotqadatasetdiverseexplainable, press2023measuringnarrowingcompositionalitygap, krishna2024factfetchreasonunified] where LLMs often perform well using only internal knowledge, we introduce WebPuzzle, a dataset designed to evaluate LLMs’ ability to locate and reason over noisy, scattered information on the open web. Figure 2 illustrates our data synthesis and curation processes.
Candidate Data Generation We collect candidate data from Wiki-corpus and real-user queries with retrieved webpages from our deployed smart assistant service. Our generation involves two approaches: (1) Cross-page question generation, where an LLM extracts facts from web pages to generate “inverted” questions, answers and checklists [multihoprag, krishna2024factfetchreasonunified]—applied only to open web pages as Wiki-corpus tends to produce overly-simple questions; and (2) Riddle creation, where the LLM selects distinctive entity attributes and applies obfuscation or generalization to create challenging problems, with original entities as labels. Examples appear in Appendix D.1. More quality assurance protocols are shown in Appendix D.3.
Difficulty Assessment To ensure stable RL training with consistent reward signals, we tag each problem’s difficulty level, enabling a data mixture strategy that prevents all-zero rewards which could lead to training collapse. For each problem, we test DeepSeek-R1 four times, using the number of correct answers to determine difficulty. The formal definition appears in Appendix D.4, with the statistics of the dataset presented in Table 7 and tagging workflow in Appendix E.6.
Test Set Annotation Unlike the training set which used LLM labeling, our test set was manually annotated by 5 human experts using an open-web search engine. From 500 seed samples, experts followed the principles in Appendix D.2 to ensure meaningful evaluation of LLMs’ information-seeking behaviors. Through iterative annotation, we finalized 275 samples for testing.